(14 over b-6) = 1/2 + (6 over b-8)
solve for b

Answer by RAY100(1637)

14/ (b-6) = 1/2 + 6/(b-8)
.
14 / (b-6) -1/2 -6/ (b-8) = 0
.
mult by 2(b-6)(b-8)
.
2*14(b-8) -(b-6)(b-8) -2*6(b-6) = 0
.
28b -224 -b^2 +14b -48 -12b +72 = 0
.
-b^2 + 30b -200 = 0
.
b^2 -30b +200 = 0
.
(b-10)(b-20) =0
.
b= 10, 20
.
check
for 10,,,, 14/4 = 1/2 +6/2,,,,3.5=3.5,,,,ok
for 20,,,,14/14 =1/2 +6/12....1=1,,,,ok